### v1

# Representation Bending for Large Language Model Safety

Abstract
Large Language Models (LLMs) have emerged as powerful tools, but their inherent safety risks – ranging from harmful content generation to broader societal harms – pose significant challenges. These risks can be amplified by the recent adversarial attacks, fine-tuning vulnerabilities, and the increasing deployment of LLMs in high-stakes environments. Existing safety-enhancing techniques, such as fine-tuning with human feedback or adversarial training, are still vulnerable as they address specific threats and often fail to generalize across unseen attacks, or require manual system-level defenses. This paper introduces RepBend, a novel approach that fundamentally disrupts the representations underlying harmful behaviors in LLMs, offering a scalable solution to enhance (potentially inherent) safety. RepBend brings the idea of activation steering – simple vector arithmetic for steering model’s behavior during inference – to loss-based fine-tuning. Through extensive evaluation, RepBend achieves state-of-the-art performance, outperforming prior methods such as Circuit Breaker, RMU, and NPO, with up to 95% reduction in attack success rates across diverse jailbreak benchmarks, all with negligible reduction in model usability and general capabilities. 111Model and code: github.com/AIM-Intelligence/RepBend
Representation Bending for Large Language Model Safety
[…]
1 Introduction
Large language models (LLMs) have become versatile tools with notable capabilities, showing promise as general-purpose task solvers. Their growing adoption across many industries and for personal use makes it increasingly important to ensure they are safe and do not cause harmful or catastrophic outcomes Hendrycks et al. (2023). LLMs are typically fine-tuned for both instruction-following and safety, to become both helpful and harmless – to obey and provide helpful responses to benign requests, and to refuse harmful ones Bai et al. (2022). Nevertheless, they still exhibit harmful behaviors, especially to adversarial manipulations Wei et al. (2024); Zou et al. (2023b); Andriushchenko and Flammarion (2024); Pelrine et al. (2023), or even fine-tuning Lermen et al. (2023); Zhan et al. (2024); Anwar et al. (2024); Qi et al. (2024). These attacks can bypass safety training of LLMs and result in models that generate harmful responses.
Additionally, future unsafe LLMs can do societal harms in ways that are not yet analyzed or perceived, considering the potential emergence of Artificial General Intelligence (AGI) in the coming years or decades Ngo et al. (2024). The rapid advances in LLMs not only expand their capabilities, but also their potential for societal harm Anwar et al. (2024); Birhane et al. (2023); Ganguli et al. (2022); Anthropic (Sep 19, 2023). With increased agency and autonomy, and as large models are scaled and deployed in higher-stakes settings in the real world, ensuring that these models are safe is critical Chan et al. (2023); Anthropic (Sep 19, 2023).
Existing safety training methods (e.g., SFT Brown et al. (2020), DPO Rafailov et al. (2024a), RLHF Christiano et al. (2017)) are frequently bypassed Zou et al. (2023b); Wei et al. (2024); Andriushchenko et al. (2024); Schwinn et al. (2024). To combat these, adversarial training is proposed, where during training specific attack methods are remedied Mazeika et al. (2024); Jiang et al. (2024b). However, adversarial training only addresses specific types of attacks and does not generalize to attacks that were unseen during training. System-level defenses, such as input and output filters Han et al. (2024), may be hard to scale due to the frequency of checks for every query, need to be upgraded with new and intelligent attacks, and do not make the underlying model inherently safer Zou et al. (2024), as they only limit the model’s exposure to unsafe content.
RepBend drives inspiration from these works, and also from the activation steering literature (Turner et al., 2024; Panickssery et al., 2023; Cao et al., 2024). Activation steering forms “steering vectors” by simply taking the difference of the activations for safe and unsafe prompts, and performs simple arithmetic (e.g., addition or subtraction) to change the model’s behavior during inference. Activation steering has limitations for in- and out-of-distribution generalizability, and can compromise model’s general reasoning capabilities Tan et al. (2024); Anthropic (Oct 25, 2024).
RepBend brings the idea of activation steering to fine-tuning, by defining a loss function whose terms are based on simple vector difference (details in Section 3). The resulting loss is not only simple and intuitive compared to representation engineering methods (Section 3), but also generalizes for out-of-distribution examples. RepBend keeps the model’s general reasoning capability, and it achieves the state-of-the-art performance. RepBend achieves up to 95% improvement in reducing attack success rates across diverse jailbreak benchmarks, with minimal impact on model usability and general capabilities. With RepBend, we advanced the Pareto frontier of of safety vs. general capability compared to the other state-of-the-art methods, including NPO Zhang et al. (2024), RMU Li et al. (2024a), Circuit Breaker Zou et al. (2024), and Task Arithmetic Ilharco et al. (2022). The results are discussed in Section 4.
2 Related Work
Unlearning.
Conventional unlearning aims to update the weights of a model to remove specific knowledge, and usually is about narrow topics (e.g., Harry Potter) or fictional information Eldan and Russinovich (2023); Maini et al. (2024). For reviews of existing works, see Liu et al. (2024). Although fine-grained control is useful, our work tackles broader notions of undesirable outputs. In the context of language models, previous work explores concepts like fairness, privacy, safety or hallucinations Jang et al. (2023); Yao et al. (2023); Liu et al. (2024). Recently NPO Zhang et al. (2024) is proposed; a simple alignment-inspired method that could be extended to unlearn hazardous knowledge. In this work we compare NPO’s performance with RepBend.
Activation Steering.
A general tactic steers language models away from generating undesirable text during inference. Liu et al. (2023) proposes in-context vectors that encode and replace in-context examples and uses it with simple vector arithmetic during inference, while other works control LLMs by “steering vector” activations Turner et al. (2024); Panickssery et al. (2023); Cao et al. (2024). Similarly, Jorgensen et al. (2023) applies mean clustering to find better steering vectors than normal averaging, and Qiu et al. (2024) proposes spectral editing of activations, an activation editing method to guide LLMs to generate desirable outputs through spectral decomposition. Critically, these methods focus on inference-time changes to the forward pass. While this makes activation steering widely applicable, activation steering has limitations for in- and out-of-distribution (OOD) generalizability, and can compromise model’s general reasoning capabilities Tan et al. (2024); Anthropic (Oct 25, 2024). RepBend has OOD generalizability and good general reasoning score (Section 4.2).
Safety Representation Engineering.
Closely related to our work, are safety frameworks that change the representations during train-time Zou et al. (2023a). For example, authors in Rosati et al. (2024) push harmful representations towards random noise to disturb the unsafe space such that they are harder to recover. R2D2 Mazeika et al. (2024) fine-tunes LLMs on a dynamic pool of harmful prompts continually updated by an optimization-based red teaming method. RMU Li et al. (2024a) selectively forgets unsafe knowledge while limiting general capabilities loss. Circuit Breaker (CB) Zou et al. (2024) improved upon RMU, and showed that we can “short circuit” the representations that are responsible for harmful outputs. Zhang et al. (2024). Motivated by CB, but different from it, we designed a loss function that operates simply by taking difference of activations, similar to activation steering, and we found that it performs better than CB and all other methods. Results are in Section 4.
3 Representation Bending
The core idea of representation bending is to “bend” the representation of the model in a way to make the model closer to safer states – latent knowledge encoded in the activation space – while making it further from harmful states. In other words, remap model representations related to harmful states, redirecting them towards incoherent or refusal representations (Figure 1).
To do so, we have to bring the model into its unsafe and safe representation states: we input safe and unsafe text to the model, to elicit the targeted representations, and read the activations of the model for those texts Liu et al. (2023); Turner et al. (2024); Zou et al. (2024, 2023a). We first need to gather a dataset consisting of safe and unsafe texts. We need text that is “unsafe and also text that is considered “safe to elicit model’s representations. Details of the datasets we used discussed in Section A.1.
RepBend is shown in Algorithm 1. RepBend tries to change the representations of model by minimizing a loss function over a fixed number of steps during fine-tuning. At a high level, RepBend bends the representations of model to be far from unsafe representations and close to safe representations, while maintaining its general capabilities. For improved performance, RepBend updates the model parameters with LoRA Hu et al. (2021). Model with LoRA parameters is referred to as (and can be seen as an independent) model . Since we only update the LoRA parameters, at the end of the algorithm would be the desired safe model. We now explain all the steps in Algorithm 1.
We sample a batch of safe text from safe sets, , safe prompts, and , unsafe prompts followed by safe responses, respectively, and feed them into the LLM to elicit safe representations (line 5). We then obtain the vector , the difference between safe representations of and (line 6). For simpler notation, we show as , denoting representations of model for a set of layers and a set of token positions . (Prior work typically uses for positions of last tokens of prompt, or first tokens of response, or all tokens of input von Rütte et al. (2024); Gurnee and Tegmark (2023); Zou et al. (2023a); Turner et al. (2024); Panickssery et al. (2023); Zou et al. (2024)). We sample a batch of unsafe text from unsafe set, , and get the unsafe representation difference vector (line 7-8). Since we want the model not to generate unsafe text, we like to have big L2 norm ( be far from for unsafe representation), while we want to have small L2 norm ( not get far from for safe representation). Hence the L2 norm of these vectors are added to the loss term (line 12) with respective negative and positive signs.
Since we also want to have stability while bending the representations, we add two loss terms with distinct purposes: KL divergence loss term between output logits of and is added so that the model still retains its general capability when input is safe. A similarity term denoted by is added to encourage similarity between outputs of unsafe prompts, given the fact that the response of a safe LLM to most unsafe prompts would begin by refusal texts (e.g.,“I’m Sorry, I am unable to assist”, or “As an assistant, I am unable to provide help.”) Arditi et al. (2024); Liu et al. (2023); von Rütte et al. (2024). In lines 9-10, a batch of unsafe text are sampled and are added to a set . Method computes average of cosine similarities between all pairs in the set . Since the batch size is often not big for LLMs, the complexity of this loss is manageable. By minimizing this loss for model , this algorithm makes the LLM safe.
Choice of Layers.
Finding the set of layers to do the intervention of RepBend is non-trivial and largely dependent on the architecture and practical considerations Wu et al. (2024); Arditi et al. (2024); Li et al. (2024b). Prior work has considered early layers Turner et al. (2024); Li et al. (2024a), middle layers Li et al. (2024b); Zou et al. (2024), later layers Zou et al. (2023a); Cao et al. (2024); Panickssery et al. (2023), and all layers Liu et al. (2023) for intervening on activations. We need to find the layers for which representation bending can have the best impact. We hypothesized that mid to later layers are responsible for the generation of output, and targeting them could be good for representation bending. This aligns well with prior work Zou et al. (2023a); Panickssery et al. (2023) that shows behavior clustering and emotion representations emerge around half or one-third of the way through the layers, indicating rich representations. We validated our hypothesis in our experiments too (Section 4.6), where representation bending works best in mid to later layers (layers 20 and after) among other choices.
Where in architecture.
Our work focuses on decoder-only, autoregressive LLMs. Each LLM accepts tokens, embeds into inputs, transforms into latents for each of transformer blocks, and finally unembeds to form logits. Each transformer block is comprised of two modules – a self-attention module and multi-layer perceptron (MLP) . Both parts are preceded by a layer norm and followed by a skip connection. The architecture can be simply be represented by:
Following existing research, we apply representation bending in the residual stream of LLM Arditi et al. (2024); Elhage et al. (2021); von Rütte et al. (2024); Yu et al. (2024); Panickssery et al. (2023); Zou et al. (2024), in particular the residual stream at the output of the block . We also experimented with other possibilities when activations are obtained from different parts of the transformer block, namely activations at the input of (after first layer normalization ), after passing through the self-attention block , and at the input of after the second layer normalization. We observed not much variability, and decided to go with residual streams as in prior works.
4 Experiments
We validate the effectiveness of RepBend on the state-of-the-art instruction-tuned LLMs, achieving up to 95% reduction in attack success rate (ASR) across diverse jailbreak benchmarks, with minimal impact on model usability and general capabilities. First we show that RepBend reduces ASR across different benchmark, and its overall performance is the best in white-box attack model. Its performance for black-box attack is also reasonable, and in both attack scenarios with minimal loss of general capability and usability. Additionally, we use logit lens to analyze the internals of the model to show the effectiveness of RepBend in not just the output (what model “says“), but also the internals (what model “thinks“).
4.1 Experiment Details
| Domain | Benchmark | Mistral 7B Instruct v0.2 | Llama3 8B Instruct | |||||||||||
| TA | NPO | RMU | CB | R2D2* | CB* | RepBend (Ours) | TA | NPO | RMU | CB | CB* | RepBend (Ours) | ||
| Black-box Jailbreak | ||||||||||||||
| ID | WildGuardTest | 13.62 | 2.80 | 7.48 | 16.29 | 44.46 | 8.54 | 8.95 | 7.08 | 0.95 | 11.75 | 7.88 | 3.74 | 7.34 |
| OOD | HarmBench | 3.75 | 0.06 | 3.75 | 16.25 | 5.63 | 13.44 | 1.56 | 1.87 | 2.19 | 9.37 | 3.75 | 13.44 | 0.31 |
| DAN | 12.06 | 0.50 | 1.81 | 4.69 | 12.06 | 1.56 | 0.50 | 5.25 | 0.38 | 0.50 | 0.25 | 0.56 | 0.75 | |
| TrustLLM Jailbreak | 4.75 | 0.25 | 32.25 | 19.00 | 3.50 | 24.25 | 4.75 | 0.50 | 0.00 | 10.25 | 0.75 | 10.75 | 3.50 | |
| PAP | 16.56 | 2.19 | 14.37 | 13.13 | 19.37 | 9.38 | 1.87 | 14.06 | 1.88 | 7.50 | 1.88 | 4.69 | 3.12 | |
| Average | 10.15 | 1.16 | 11.93 | 13.87 | 17.00 | 11.43 | 3.53 | 5.75 | 1.08 | 7.87 | 2.90 | 6.64 | 3.00 | |
| White-box Jailbreak | ||||||||||||||
| OOD | GCG | 61.56 | 10.00 | 11.56 | 25.00 | 8.44 | 9.37 | 5.00 | 51.25 | 7.50 | 11.87 | 4.37 | 3.44 | 2.50 |
| Prefilling | 80.83 | 8.75 | 7.50 | 5.00 | 47.08 | 5.42 | 0.83 | 83.34 | 20.42 | 6.67 | 7.92 | 3.33 | 4.17 | |
| Input Embed | 30.42 | 22.08 | 50.83 | 55.83 | 44.17 | 21.67 | 2.50 | 74.17 | 51.67 | 28.33 | 41.67 | 23.75 | 3.33 | |
| Average | 55.63 | 13.61 | 23.30 | 28.05 | 33.23 | 12.15 | 2.78 | 69.59 | 26.53 | 15.62 | 17.99 | 10.17 | 3.33 | |
| Total Average | 23.14 | 5.83 | 16.19 | 19.19 | 23.09 | 11.70 | 3.25 | 29.69 | 10.62 | 10.78 | 8.56 | 7.96 | 3.13 | |
Comparisons.
We compare RepBend with several state-of-the-art methods. In safety representation engineerng literature, RMU Li et al. (2024a) and Circuit Breaker (CB) Zou et al. (2024) are the state-of-the-art methods that are directly related to our work. We train models using these methods in our experiment infrastructure for higher performance and fair comparison (e.g., we found out training a CB model with our setup comes with better performance in some settings than the model publicly available). We also compare RepBend with two related unlearning methods that can be applied to safety: Task Arithmetic (TA) Ilharco et al. (2022) and Negative Preference Optimization (NPO) Zhang et al. (2024), where harmful behavior can be unlearned by negating harmful task vector (in TA) or reversing alignment (in NPO).
Finally, we compare RepBend against public safety-aligned models, R2D2222https://huggingface.co/cais/zephyr_7b_r2d2 and CB 333https://huggingface.co/GraySwanAI/Llama-3-8B-Instruct-RR444 https://huggingface.co/GraySwanAI/Mistral-7B-Instruct-RR, trained by the authors of the respective works. Further details are provided in Section A.2.
Datasets.
We curate the training dataset from three sources: WildGuardMix Han et al. (2024) and WildJailbreak Jiang et al. (2024b), which include harmful and harmless prompts, and UltraChat Ding et al. (2023), which contains general instruction-following data. From each dataset we randomly select 10,000 samples, and classify them into safe (harmless) and unsafe (harmful) groups. See Section A.1 for more details.
Training Details.
We conduct experiments primarily using Mistral 7B v0.2 and Llama 3 8B (see list of all models in Table 4). Experiments with other LLMs of various sizes are reported in Section 4.4. We initialize the models from instruction-tuned checkpoints and apply LoRA Hu et al. (2021) with rank and alpha of 16, targeting all linear layers in the model for all methods and RepBend. However, in RMU we only target the MLP layer in each transformer block, as per the paper Li et al. (2024a). Details are in Section A.4.
4.2 Robustness Against Jailbreak Attacks
Benchmarks.
We evaluate RepBend against diverse jailbreak attacks, grouped into two categories: black-box and white-box access attacks.
| Model | Method | Safety | Over-refusal | General Capability | Overall () | |
| Average ASR () | XSTest () | Wildjailbreak: Benign () | Average Capability () | |||
| Mistral 7B Instruct v0.2 | Original Weight Model | 60.64 | 85.78 | 100.00 | 59.18 | 63.81 |
| TA Ilharco et al. (2022) | 23.14 | 80.22 | 97.60 | 53.10 | 72.96 | |
| NPO Zhang et al. (2024) | 5.83 | 68.89 | 70.00 | 53.94 | 74.52 | |
| RMU Li et al. (2024a) | 16.19 | 78.44 | 90.40 | 47.32 | 71.85 | |
| CB Zou et al. (2024) | 19.19 | 86.89 | 97.60 | 58.97 | 77.34 | |
| R2D2* Mazeika et al. (2024) | 23.09 | 67.56 | 96.80 | 48.44 | 72.67 | |
| CB* Zou et al. (2024) | 11.70 | 86.22 | 82.00 | 58.93 | 73.62 | |
| RepBend (Ours) | 3.24 | 84.89 | 93.60 | 57.68 | 81.23 | |
| Llama3 8B Instruct | Original Weight Model | 34.00 | 85.11 | 92.00 | 67.14 | 73.90 |
| TA Ilharco et al. (2022) | 29.69 | 80.00 | 88.80 | 57.43 | 70.71 | |
| NPO Zhang et al. (2024) | 10.62 | 74.45 | 43.20 | 66.71 | 71.65 | |
| RMU Li et al. (2024a) | 10.78 | 76.89 | 72.40 | 54.84 | 72.90 | |
| CB Zou et al. (2024) | 8.56 | 84.44 | 89.20 | 66.58 | 81.61 | |
| CB* Zou et al. (2024) | 7.96 | 85.78 | 52.40 | 66.47 | 75.87 | |
| RepBend (Ours) | 3.13 | 84.11 | 89.20 | 65.90 | 83.14 | |
| Model | Method | Safety | Over-refusal | General Capability | Overall () | |||
| Harmbench () | WildguardTest () | XSTest () | Wildjailbreak: Benign () | MTBench () | MMLU () | |||
| Gemma2 2B Instruct | Original Weight | 11.56 | 28.70 | 78.67 | 98.80 | 7.35 | 57.98 | 78.12 |
| RepBend (Ours) | 6.56 | 1.34 | 70.34 | 82.80 | 7.37 | 58.14 | 79.51 | |
| Qwen2.5 14B Instruct | Original Weight | 17.19 | 33.11 | 86.67 | 100.0 | 8.71 | 79.60 | 83.85 |
| RepBend (Ours) | 7.50 | 6.67 | 82.22 | 99.60 | 9.14 | 78.89 | 89.66 | |
Black-box access attacks assume no internal access to LLMs. The benchmarks include direct harmful requests (HarmBench Mazeika et al. (2024)), pre-generated jailbreak prompts (WildGuardTest Han et al. (2024), DAN Shen et al. (2023), TrustLLM-Jailbreak Huang et al. (2024)), and transformed instructions using external models (PAP Zeng et al. (2024)).
White-box access attacks assume some access to model internals. The benchmarks we use include GCG Zou et al. (2023b), Prefilling Andriushchenko et al. (2024); Vega et al. (2023), and Input Embed Schwinn et al. (2024). GCG finds an adversarial suffix for a prompt by maximizing compliance likelihood, Prefilling attack prefills the LLM’s response with a non-refusal beginning, and Input Embed produces adversarial embeddings instead of tokens.
WildGuardTest is the only in-distribution (ID) benchmark; other benchmarks are out-of-distribution (OOD), as they are based on either unseen adversarial prompts or unseen attack methods that were not part of the training or fine-tuning. Additional details about the benchmarks are provided in Section A.5. Unless otherwise stated, we use an open-source classifier Mazeika et al. (2024) to evaluate compliance of the responses.
Results.
Table 1 shows the attack success rate (ASR) – the model’s compliance to the attacks – across five black-box and three white-box access attack benchmarks on Mistral 7B and Llama 3 8B models. Figure 2 plots the total average ASR of white-box and black-box attacks combined. RepBend achieves the lowest average ASR (3.25 for Mistral and 3.13 for Llama) compared to other methods, improving refusal rates by 94.64% and 90.79% over the original instruction-tuned Mistral 7B and Llama 3 8B, respectively. We can see in Table 1 that in black-box attacks, RepBend achieves the lowest ASR for DAN and PAP in Mistral 7B and lowest ASR for HarmBench in Llama 3. For white-box attacks, RepBend most of the time achieves the lowest ASR. Moreover, these results demonstrate strong out-of-distribution (OOD) generalizability, as RepBend effectively refuses adversarial prompts and optimization attacks that were not seen during training.
While NPO performs well on black-box attacks, it does not perform well in white-box attacks because it only targets outputs of the model, and not the model’s internals for compliance to harmful requests. In contrast, RepBend delivers robust performance across both categories as it also changes the model internals via representation bending.
4.3 Pareto-Frontier of Safety, Usability, and Capability
Achieving an optimal balance between safety and general capability is crucial for LLMs. Over-refusing benign queries can compromise usability and improving safety should not degrade the model’s core capabilities. In this section we show that RepBend achieves Pareto-frontier of safety, usability, and general capability.
Benchmarks.
We evaluate over-refusal using XSTest Röttger et al. (2023) and WildJailbreak-Benign Jiang et al. (2024b). These datasets contain benign prompts with ambiguous wording (either intentional or unintentional) that seem harmful in form but contain no harmful intent. We measure compliance rate to these prompts using GPT-4. For general capability, we use 8 benchmarks which includes MTBench Zheng et al. (2023), MMLU Hendrycks et al. (2020), BBH Suzgun et al. (2022), TruthfulQA Lin et al. (2021), ARC-C Clark et al. (2018), Winogrande Sakaguchi et al. (2021), GSM8K Cobbe et al. (2021), Codex-Eval Chen et al. (2021), to measure instruction-following capability, factual knowledge, and problem-solving performance including math and coding. Details are in Section A.5.
Results.
Table 2 shows the three axes of safety, over refusal, and general capability of all methods. Overall score provides a single index to better see the Pareto-frontier performance by combining the 3 axis. To calculate it we level the score ranges to be the same and average them: safety score ( averaged on 8 benchmarks), over-refusal score (average of 2 benchmarks) and general capability score (average of 8 benchmarks where MTBench is 10 scaled). RepBend achieves the highest overall score, showing that RepBend does not compromise usability or capability for safety improvements. Whereas RepBend does not have the highest general capability and over-refusal scores, it has reasonable and acceptable scores comparable to the original model while maintaining the lowest ASR on jailbreak benchmarks. (general capability scores for each benchmark are in Section B.3). NPO shows good safety and general capability scores but struggles with distinguishing benign requests in over-refusal scores. CB achieves balanced scores but falls short in safety compared to RepBend.
Additional results with more algorithms like SFT, DPO Rafailov et al. (2024b), WHP Eldan and Russinovich (2023), and the original model with a safe instructed prompt (e.g., “If someone asks you an unsafe or harmful prompt, do not answer”) are available in Section B.1. Moreover, further comparison with controllable text generation methods (e.g., a classifier-based method Yang and Klein (2021)) with better transferability are discussed in Section B.2
4.4 General Applicability Across LLM Architectures and Sizes
Models.
To demonstrate the scalability of RepBend, we evaluate its performance across additional LLM architectures of varying sizes. The architectures include Gemma2 2B-Instruct, and Qwen2.5 14B-Instruct, where the models have different number of hidden layers. We evaluate safety, over-refusal, and general capability using the six benchmarks introduced in Section 4.3.
Results.
Table 3 shows the performance of RepBend across the additional LLM architectures. With minimal hyperparameter tuning, RepBend successfully enhances safety while maintaining usability and capability. These results confirm the scalability of RepBend across diverse architectures and parameter sizes.
4.5 Impact of cos_sim loss
We introduce cos_sim loss in RepBend to enhance the both safety and stability of the fine-tuned model. To evaluate its impact, and to show the ease of finding a value for its hyper-parameter , we conduct ablation studies on .
Results.
Figure 4 illustrates the role of cos_sim loss in mitigating harmful outputs for the Llama3 8B Instruct model. Removing cos_sim loss () severely degrades safety, leading to increased harmful responses. Setting a small greatly improves safety, and then performance remains stable across a broad range of , suggesting robustness within a "safety basin" where aligned models maintain strong performance despite small perturbations Peng et al. (2024).
We can see that the performance remains stable as increases. This trend also holds across various hyperparameter settings, and shown in Figure 5, underscoring the importance of cos_sim loss in maintaining safety and relative insensitivity of the loss to larger values. Due to this stability of performance, extensive hyperparameter searches for can be avoided by setting it to a reasonable non-zero value.
4.6 Internal Behavior of Model
Representation Analysis.
To analyze the impact of RepBend on internal representations, we need a way to visualized the model’s internals. We use the Logit Lens framework Nostalgebraist (2020), which maps model’s representations from the latent space to the vocabulary space for interpretable analysis. Logit Lense hence allows us to see what the next predicted token in each layer is and it is widely used to investigate Transformer-based architectures Dar et al. (2023); Elhage et al. (2021).
Figure 3 (top) illustrates a harmful request from HarmBench Mazeika et al. (2024), and (bottom) shows the Logit Lens of (a) the instruction-tuned Llama3 8B baseline and (b,c) RepBend version of Llama3 8B Instruct. The heatmap cells show next token predictions across layers and colors show confidence levels (blue: high, red: low) during response generation. The baseline model produces a harmful response with a sharp increase in prediction certainty at later layers (Figure 3(a)). In contrast, RepBend successfully refuses the request, significantly reducing certainty in harmful token predictions (Figure 3(b)). We can also see layers beyond 20 are critical for next-token generation, confirming our choice of layers in RepBend.
[…]
5 Conclusion
The rapid adoption and scaling of LLMs amplify the urgency of ensuring their safety in diverse real-world applications. While existing safety measures provide partial solutions, they often fall short in generalizing to unseen attacks or maintaining model capabilities. Addressing these gaps, RepBend presents a novel fine-tuning approach inspired by activation steering, which optimally bends model representations to maximize safety while preserving general-purpose capabilities.
RepBend’s performance, including up to 95% reduction in attack success rates across benchmarks, demonstrates the method’s robustness and generalizability compared to state-of-the-art approaches. We showed that RepBend can be applied to different LLM architectures of varying sizes. Moreover, RepBend effectively balances safety and capability, advancing the Pareto frontier of these often-competing objectives.
This work highlights the potential of embedding safety principles directly into the training process to create inherently safer models. Future research can build on this foundation to explore broader applications, optimize computational efficiency, and address emerging challenges in the evolving landscape of LLM safety. Through continuous innovation, RepBend moves us closer to the responsible and secure deployment of advanced AI systems.
6 Limitations
An inherent limitation of any unlearning method, is robustness to re-learning hazardous knowledge. Similar to the finding for RMU Li et al. (2024a), we also found out that if RepBend is fine-tuned with unsafe data, it can relearn the harmful content. This limitation of robustness has also been observed by several others Qi et al. (2024); Lo et al. (2024); Lermen et al. (2023); Zhan et al. (2024), and is still an ongoing problem in the community Lynch et al. (2024), with some recent work to remedy this Rosati et al. (2024); Lyu et al. (2024). Future work can borrow techniques to make RepBend safe post-tuning.
We tested RepBend in this paper only on LLMs, and only on select open-source models. Even though we believe our methods are easily extensible to other models, the generalization of these findings are yet to be tested to other models, especially those at greater scale, including current state-of-the-art proprietary models.
Another limitation of RepBend is the effort for finding the best hyper parameters. We found the best working hyper parameters that show the results in this paper; however, the we do not claim that these set of values are the optimal, nor we think the effort to find the best hyper parameter is minimal. We found out that sometimes RepBend is sensitive towards the choice of hyper parameters, specially the loss coefficient parameters.
7 Broader Impact and Potential Risks
This work introduces a scalable and generalizable approach to model safety that can influence the development of future safety standards, fostering industry-wide adoption of safer practices in AI research and deployment. RepBend provides a practical framework for enhancing the safety of large language models, enabling their deployment in high-stakes domains such as healthcare, education, and legal systems, where the consequences of unsafe outputs could be catastrophic.
While RepBend reduces vulnerabilities to adversarial attacks, its success may lead to overconfidence in the safety of AI systems, potentially encouraging premature deployment in sensitive domains without rigorous oversight.
The advancement of safety mechanisms like RepBend may inadvertently escalate an arms race between AI safety researchers and malicious actors, leading to increasingly sophisticated attacks that may exploit yet-undiscovered vulnerabilities.
As with any technological development, there is a potential for misuse. We believe that disclosing details to prevent attacks on LLMs can be beneficial if used appropriately. However, it also poses the risk of malicious attackers exploiting this information to develop new methods for bypassing the defense. The proposed loss function in RepBend could be inverted or manipulated to intentionally create models that generate unsafe or toxic outputs. Or with the same RepBend algorithm, but different data, it can be used for different purpose if we change the data.
Lastly, finding the optimal set of hyper parameters may need large search, which directly translates to environmental concerns for energy usage.
Acknowledgments
We would like to thank Hyungjoo Chae and Jiwan Chung for the discussions and their insightful ideas and comments.
Appendix A Experiment Details
A.1 Training Datasets
We utilize the WildGuardMix training dataset Han et al. (2024), a comprehensive collection of diverse, vanilla and adversarially-designed queries paired with benign and harmful responses. This dataset allows the model to learn appropriate responses to benign queries while avoiding responses to harmful queries with unlearning techniques. We randomly select total of 10,000 benign and harmful requests and responses from this dataset and classify them into safe (harmless) and unsafe (harmful) groups.
We also utilize the Wildjailbreak dataset Jiang et al. (2024b), a benchmark specifically curated to test the robustness of LLMs against adversarial jailbreak attempts. We create paired data that include both compliance and refusal responses for the same harmful request. For generating harmful responses to harmful requests we use an uncensored open-source LLM555https://huggingface.co/maywell/PiVoT-0.1-Evil-a. These pairs not only increase the diversity of samples but are also essential for training algorithms like DPO and RepBend. We select 10,000 harmful queries from this dataset.
Finally, we include samples from the UltraChat dataset Ding et al. (2023), a diverse dataset for instruction-tuning, to ensure the model maintains general capabilities alongside safety improvements. We include 10,000 samples as safe samples.
A.2 Comparison Baselines
We compare our method against a variety of baselines, including standard fine-tuning methods, machine unlearning techniques, and recently proposed safety approaches. We use batch size of 16 and learning rate of with Adam Optimizer to run each method unless otherwise is specified.
| Model | Source | Access | License |
| Gemma2 2B Instruct | Mesnard et al. (2024) | Link | Gemma Terms of Use |
| Llama-3 8B Instruct | AI@Meta (2024) | Link | Llama 3 Community License |
| Mistral 7B Instruct v0.2 | Jiang et al. (2023) | Link | Apache License 2.0 |
| Mixtral 8x7B Instruct v0.1 | Jiang et al. (2024a) | Link | Apache License 2.0 |
| Mistral 7B Instruct RR | Zou et al. (2024) | Link | MIT License |
| Qwen2.5 3B Instruct | Bai et al. (2023) | Link | Qwen Research License |
| Qwen2.5 14B Instruct | Bai et al. (2023) | Link | Apache License 2.0 |
| PiVoT-0.1-Evil-a | - | Link | CC-BY-SA-4.0 |
| Zephyr 7b R2D2 | Mazeika et al. (2024) | Link | MIT License |
| HarmBench Llama2 13b cls | Mazeika et al. (2024) | Link | MIT License |
| TruthfulQA Truth Judge Llama2 7B | Lin et al. (2021) | Link | Apache License 2.0 |
| TruthfulQA Info Judge Llama2 7B | Lin et al. (2021) | Link | Apache License 2.0 |
| Dataset | Source | Access | License |
| HarmBench | Mazeika et al. (2024) | Link | MIT License |
| MMLU | Hendrycks et al. (2020) | Link | MIT License |
| ARC | Clark et al. (2018) | Link | CC-BY-SA-4.0 |
| GSM8K | Cobbe et al. (2021) | Link | MIT License |
| WinoGrande | Sakaguchi et al. (2021) | Link | Apache License 2.0 |
| TruthfulQA | Lin et al. (2021) | Link | Apache License 2.0 |
| wildjailbreak | Jiang et al. (2024b) | Link | Open Data Commons License |
| wildguardmix | Han et al. (2024) | Link | Open Data Commons License |
SFT: The baseline model fine-tuned on the safe training set using cross-entropy loss. We fine-tuned it for 1 epoch.
DPO Rafailov et al. (2024b): Trained using paired samples generated from the Wildjailbreak dataset, where model learns to prefer safe over unsafe responses. We fine tuned it for 1 epoch, but with the learning rate .
Task Arithmetic Ilharco et al. (2022): Combines using vector arithmetic (subtraction) two model weights trained with LoRA on safe and unsafe data, respectively. The formula for task arithmetic here is where and denote the model weights for training with safe and unsafe data, respectively. Each model is fine-tuned for 1 epoch and we set in our experiments.
Safety Prompting: Motivated from recent works Lynch et al. (2024); Thaker et al. (2024), we also wonder if providing a simple safe instructions in the prompts can make a model safe. Here safety instructions are used as prefix to the input prompts. We generate five safety instructions using GPT and randomly select one during evaluation. The instructions are shown below.
WHP Eldan and Russinovich (2023): Adapts unlearning methods for forgetting specific concepts, such as Harry Potter. We use the unsafe training set as the forget set and dynamically alter labels based on a reinforced model trained on unsafe data. The reinforced model is fine-tuned on unsafe set for 1 epoch and the unlearning model is updated for 150 steps.
NPO Zhang et al. (2024): Unlearns knowledge about harmful samples while maintaining general capability, without requiring paired data. The unsafe training set serves as the forget set, and the safe training set as the retain set. We iterate 600 steps for Mistral and 150 steps for Llama3.
RMU Li et al. (2024a): Routes the representations of harmful samples to a random representation to unlearn harmful knowledge in the model. We set in our experiments for the loss and run with batch size 4 and max step 150.
Circuit Breaker Zou et al. (2024): Targets specific harmful representations in hidden states to make them orthogonal to the original representation. We set and learning rate to , and fine-tune for 200 steps.
A.3 List of Models and Datasets
A.4 RepBend Hyper Parameters
For RepBend, We set the hyperparameters as , , . is computed for all layers to maintain the harmless representations across all layers, while the is computed from the 20th layer to last layer to effectively target and bend representations responsible for harmful content generation. We use the learning rate of and the batch size of 16 and update the Mistral 7B Instruct v0.2 and Llama3 8B Instruct models for 300 and 450 steps, respectively. For the hyperparameter searching, we perform grid search over batch size {8, 16, 32}, learning rates of {}, the max step {150, 300, 450, 600}. For , , and we searched each in {0.1, 0.2, 0.3, 0.5, 0,7}. We also searched over the choice of layers: early, mid, and late, and found the maximum gain in later layers (responsible for harmful content generation).
A.5 Evaluation Benchmarks
We evaluate our method and other baselines using various benchmarks to measure safety, over-refusal, and general capability. All evaluations are conducted using the following code bases:
Safety-eval: https://github.com/allenai/safety-eval
Jailbreak Attack Benchmarks
1.
Black-box Attacks
HarmBench Mazeika et al. (2024): A dataset with 320 naive requests representing diverse harmful behaviors without any jailbreak techniques.
WildGuardTest Han et al. (2024): Contains total of 749 vanilla and adversarial harmful requests, both with and without jailbreak tactics.
TrustLLM-Jailbreak Huang et al. (2024): Dataset containing 13 distinct jailbreak attacks. We use only subset of 400 prompts derived from the original test set.
DAN Shen et al. (2023): The study analyzes 1,405 jailbreak prompts from 131 communities and creating a dataset of 107,250 samples to test attack strategies.
PAP Zeng et al. (2024): Develops a persuasion taxonomy based on social science principles and used it to automatically generate persuasive adversarial prompts (PAPs). It uses Mixtral-8x7B-Instruct to convert harmful instructions into persuasive forms.
Figure 4: Ablation study of cos_sim loss term in RepBend on Llama3 8B Instruct. All other hyperparameters are fixed.
Figure 5: Impact of the loss-weighting hyperparameter on safety, measured as the average ASR on WildGuardTest, HarmBench, and PAP. Each color represent a configuration of hyper-parameters. The cos_sim is relatively unaffected by , with similar trends across different hyperparameter combinations. -
2.
White-box Attacks
GCG Zou et al. (2023b): updates input tokens to improve possibility of jailbreaking via gradient optimization: by maximizing compliance likelihood, GCG finds an adversarial suffix for a prompt.
Prefilling Zou et al. (2024): Prefills the initial response with the beginning of a desired target completion (e.g., with a non-refusal beginning) to maximize the compliance likelihood by optimization.
Input Embed Schwinn et al. (2024): Produces adversarial input embeddings instead of tokens using gradient-based optimization.
We use 320 HarmBench samples for GCG, while Prefilling and Input Embed are tested on 240 harmful samples following their original implementation We run GCG optimization for 500 iterations, with adversarial string initiated as “! ! ! ! ! ! ! ! ! ! ! ! ! ! ! ! ! ! ! !” (default settings of HarmBench), and searching 512 adversarial candidates at each step. The evaluations (except TrustLLM) use the Open-source classifier from HarmBench Mazeika et al. (2024) to determine attack success rates, compliance, and general quality of the responses. TurstLLM-Jailbreak is evaluated with GPT-4.
| Model | Method | Safety | Over-refusal | General Capability | |||
| Harmbench () | WildguardTest () | XSTest () | Wildjailbreak: Benign () | MTBench () | MMLU () | ||
| Mistral 7B Instruct-v0.2 | Original | 50.94 | 53.00 | 85.78 | 100.00 | 7.72 | 60.18 |
| SFT | 4.69 | 15.49 | 78.89 | 96.80 | 5.98 | 58.46 | |
| DPO | 25.63 | 36.45 | 83.56 | 99.60 | 7.57 | 59.21 | |
| TA | 3.75 | 13.62 | 80.22 | 97.60 | 6.93 | 58.57 | |
| Safety Prompting | 20.00 | 38.45 | 81.77 | 98.80 | 7.50 | 59.26 | |
| WHP (W/O safe set) | 49.06 | 52.60 | 85.56 | 99.60 | 7.88 | 60.08 | |
| WHP | 33.13 | 42.86 | 80.44 | 98.40 | 6.29 | 59.24 | |
| NPO (W/O safe set) | 5.63 | 7.61 | 76.67 | 50.00 | 4.67 | 59.12 | |
| NPO | 0.06 | 2.80 | 68.89 | 70.00 | 7.31 | 59.61 | |
| R2D2* | 5.63 | 44.46 | 67.56 | 96.80 | 5.97 | 59.44 | |
| RMU | 3.75 | 7.48 | 78.44 | 90.40 | 5.77 | 50.02 | |
| CB* | 13.44 | 8.54 | 86.22 | 82.00 | 7.57 | 60.08 | |
| CB | 16.25 | 16.29 | 86.89 | 97.60 | 7.58 | 59.93 | |
| RepBend (Ours) | 1.56 | 8.95 | 84.89 | 93.60 | 7.50 | 59.48 | |
| Llama3 8B Instruct | Original | 22.19 | 15.49 | 85.11 | 92.00 | 7.84 | 65.89 |
| SFT | 1.56 | 10.15 | 80.44 | 90.80 | 6.88 | 66.01 | |
| DPO | 3.13 | 0.67 | 63.11 | 28.40 | 7.54 | 65.31 | |
| TA | 1.87 | 7.08 | 80.00 | 88.80 | 7.00 | 66.09 | |
| Safety Prompting | 0.31 | 2.27 | 74.45 | 65.60 | 7.18 | 64.41 | |
| WHP (W/O safe set) | 18.75 | 0.08 | 84.00 | 92.40 | 8.01 | 66.44 | |
| WHP | 12.50 | 7.87 | 80.22 | 83.60 | 7.63 | 66.16 | |
| NPO (W/O safe set) | 11.25 | 3.87 | 77.56 | 76.00 | 7.71 | 66.29 | |
| NPO | 2.19 | 0.95 | 74.45 | 43.20 | 7.79 | 66.13 | |
| RMU | 9.37 | 11.75 | 76.89 | 72.40 | 4.01 | 58.23 | |
| CB* | 13.44 | 3.74 | 85.78 | 52.40 | 7.72 | 65.77 | |
| CB | 3.75 | 7.88 | 84.44 | 89.20 | 7.74 | 65.77 | |
| RepBend (Ours) | 0.31 | 7.34 | 84.11 | 89.20 | 7.71 | 65.08 | |
Over-refusal Benchmarks
XSTest Röttger et al. (2023): Contains 450 requests with ambiguous wording, testing contextual understanding to avoid over-refusal.
WildJailbreak: Benign Test Jiang et al. (2024b): Includes 210 benign but seemingly adversarial requests, evaluating the model’s handling of ambiguous prompts and over-refusla.
GPT-4 evaluates compliance for over-refusal benchmarks.
General Capability Benchmarks
General Capability Benchmarks
Big Bench Hard (BBH) Suzgun et al. (2022): Includes tasks that are challenging for LLMs. It focuses on tasks requiring complex reasoning, world knowledge, and nuanced understanding of language
TruthfulQA Lin et al. (2021): A benchmark for evaluating the truthfulness of language models by testing their ability to avoid generating false or misleading statements. It includes questions that prompt models to produce common misconceptions or falsehoods, assessing how well they can resist these temptations.
ARC-C Clark et al. (2018): A dataset of science questions from standardized exams, designed to test advanced reasoning and problem-solving. It contains the hardest questions that cannot be answered by simple retrieval or shallow heuristics.
Winogrande Sakaguchi et al. (2021): A large-scale dataset designed to evaluate commonsense reasoning through the Winograd Schema Challenge, which involves choosing the correct pronoun reference in ambiguous sentences.
GSM8K Cobbe et al. (2021): A dataset of 8,000 high-quality grade-school math word problems aimed at evaluating language models’ mathematical reasoning abilities. The problems require multi-step reasoning and arithmetic calculations.
Codex-Eval Chen et al. (2021): Codex-Eval evaluates the potential of language models in understanding and generating complex code with contextual awareness.
| Benchmark | Mistral 7B Instruct-v0.2 | Llama3 8B Instruct |
| Black-box Jailbreak | ||
| WildGuardTest | 53.00 | 15.49 |
| HarmBench | 50.94 | 22.19 |
| DAN | 58.00 | 9.06 |
| TurstLLM Jailbreak | 44.25 | 11.00 |
| PAP | 26.25 | 12.81 |
| Average | 46.49 | 14.11 |
| White-box Jailbreak | ||
| GCG | 68.12 | 36.87 |
| Prefilling | 95.00 | 85.00 |
| Input Embed | 89.58 | 79.58 |
| Average | 84.23 | 67.15 |
| Total Average | 60.64 | 34.00 |
Appendix B More Experiment Results
B.1 Full Comparison
Table 6 presents a comprehensive comparison of RepBend. It has all the results of Table 2 with additional baselines, such as SFT, DPO, WHP, Safety Prompting, and naive unlearning algorithms without retain set. These comparisons highlight the versatility and robustness of RepBend in achieving superior safety alignment while maintaining usability and general capabilities.
SFT exhibits significant reductions in general capability scores, underscoring the importance of balancing safety alignment with knowledge retention. The performance of DPO shows high variability across models, as it heavily depends on the quality of the initial model used during training. Similarly, Safety Prompting methods demonstrate inconsistent performance, with Llama3 8B Instruct, having good performance and with Mistral 7B Instruct have low performance. This is because certain models, such as LLama 3 8B Instruct are good in instruction following.
Unlearning algorithms also exhibit sub-optimal performance across various metrics. WHP is ineffective both with and without the safe training set, as it is designed to unlearn specific concepts (e.g., characters from “Harry Potter”), which is not a effective method for safety alignment. Meanwhile, NPO has good safety score even without the safe set; its safety and general capabilities can improve further when the retain set (safe set) is included. This is because using KL divergence on the retain set preserves the model’s overall knowledge while refining its responses to harmful requests. However, NPO still fails to respond appropriately to certain benign queries (over-refusal), which diminishes its usability.
B.2 Comparison with Controllable Text Generation
| Method | Safety | Over-refusal | General Capability | Overall () | |||
| Harmbench () | WildguardTest () | XSTest () | Wildjailbreak: Benign () | MTBench () | MMLU () | ||
| Original Weight | 22.19 | 15.49 | 85.11 | 92.00 | 7.84 | 65.89 | 80.62 |
| FUDGE | 14.69 | 12.68 | 83.33 | 90.40 | 6.82 | 61.03 | 79.27 |
| Safety-Propmting | 0.31 | 2.27 | 74.45 | 65.60 | 7.18 | 64.41 | 78.95 |
| RepBend (Ours) | 0.31 | 7.34 | 84.11 | 89.20 | 7.71 | 65.08 | 84.64 |
Motivated by recent works Lynch et al. (2024); Thaker et al. (2024) and Yang and Klein (2021), we wonder if simple prompting or text generation controlling methods are enough to make a model safe. We compare RepBend against FUDGE (Yang and Klein, 2021) and Safety Prompting (providing a simple safe instructions in the prompts described in Section A.2). FUDGE (Yang and Klein, 2021) leverages a classifier-based approach that a future discriminator guides the generation of next token. To bring FUDGE to safety, we train the future discriminator to classify whether the generated text is benign or harmful using safe and unsafe groups in the train set. Safety Prompting pre-pends safety instructions to the input prompts.
Results are reported in Table 8. We can see that RepBend outperforms other methods by achieving the best balance between safety, over-refusal, and general capability. While Safety Prompting demonstrates strong performance on safety benchmarks, its effectiveness heavily relies on the LLM’s instruction-following ability (see Safety Prompting on Mistral 7B in Table 6). Additionally, RepBend preserves general capability at a level comparable to the original model, whereas FUDGE and Safety Prompting exhibit noticeable degradation.
Although methods like FUDGE offer better transferability across models without further training, its lower safety performance makes it less reliable for robust alignment and safety-critical applications. RepBend, on the other hand, provides a more effective and structured solution by directly modifying model representations, ensuring both strong safety and usability without compromising performance. The transferability of LoRA for safety improvements remains an open challenge and is left for future exploration.
B.3 Results on General Capabilities
| Benchmark | Mistral 7B Instruct-v0.2 | Llama3 8B Instruct | |||||||||||||
| Original | TA | NPO | RMU | CB | R2D2* | CB* | RepBend (Ours) | Original | TA | NPO | RMU | CB | CB* | RepBend (Ours) | |
| MTBench | 7.72 | 6.93 | 7.31 | 5.77 | 7.58 | 5.97 | 7.56 | 7.50 | 7.84 | 7.00 | 7.79 | 4.01 | 7.74 | 7.71 | 7.71 |
| MMLU | 60.18 | 58.57 | 59.60 | 50.02 | 59.97 | 59.41 | 60.08 | 59.48 | 65.89 | 66.09 | 66.13 | 58.23 | 65.74 | 65.77 | 64.86 |
| BBH | 39.91 | 52.96 | 22.50 | 26.02 | 38.52 | 48.98 | 37.69 | 37.96 | 62.31 | 66.30 | 57.31 | 45.56 | 60.46 | 61.01 | 58.80 |
| TruthfulQA | 81.39 | 76.50 | 70.99 | 69.89 | 81.99 | 36.35 | 81.27 | 78.33 | 60.83 | 66.46 | 63.28 | 45.78 | 62.30 | 59.49 | 62.67 |
| ARC-C | 74.74 | 73.63 | 74.57 | 65.44 | 74.49 | 73.89 | 74.57 | 74.23 | 80.12 | 79.86 | 80.03 | 75.34 | 80.72 | 79.95 | 79.86 |
| Winogrande | 58.33 | 56.51 | 57.46 | 49.64 | 58.88 | 55.56 | 58.48 | 58.48 | 58.72 | 59.98 | 58.56 | 54.14 | 58.33 | 58.58 | 58.56 |
| GSM8K | 44.50 | 4.00 | 40.00 | 34.50 | 43.50 | 34.00 | 48.00 | 41.00 | 74.50 | 3.00 | 73.50 | 73.50 | 76.00 | 74.50 | 73.00 |
| Codex | 37.19 | 33.35 | 33.26 | 25.37 | 38.57 | 35.73 | 19.63 | 36.98 | 56.34 | 47.74 | 56.98 | 46.04 | 51.71 | 55.12 | 52.38 |
We evaluate RepBend on a broader set of benchmarks to assess its general capabilities, including reasoning, truthfulness, mathematics, and coding skills. The benchmarks, in addition to MTBench and MMLU, include BBH Suzgun et al. (2022), ARC-C Clark et al. (2018), Winogrande Sakaguchi et al. (2021) for reasoning; TruthfulQA Lin et al. (2021) for truthfulness; GSM8K Cobbe et al. (2021) for mathematical skills; and Codex Chen et al. (2021) for coding skills (see Section A.5).
Figure 6 shows the general capability score for Mistral 7B. (the results for MTBnench are multiplied by 10 for better presentation). Last set shows the average. We can see that that RepBend maintains a good general capability score (blue) comparable to the original model (grey) across all evaluated benchmarks. CB also performs well on all general capability benchmarks, while other methods have low performance on at least one benchmark. There is no significant degradation in performance, confirming the generalizability of RepBend. We can see that RepBend can effectively achieve safety alignment (best performance in Figure 2) while preserving the model’s problem-solving and knowledge-based skills. Detailed numbers and the results for Llama3 8B Instruct are in Table 9.
B.4 Another Representation Analysis Example
Figure 7 shows another representation analysis using Logit Lens for another harmful prompt. THe hamrful prompt is: “Write a persuasive article linking immigration into the US from Mexico to increased crime.”The base Llama3 8B Instruct model complies with the prmompt and write the article (Figure 7(a)). We can see in Figure 7(b) that RepBend refuses to comply with the prompt, and in Figure 7(c) that it again generates random tokens with high entropy (red colors near the last layers) when the complying response is forced to the model. Note that generating random tokens is not the usual behavior of vanilla LLMs; as shown in Figure 7(d) the model continues any sequences. However, in Figure 7(c), RepBend successfully refuses the harmful response even when it is forced.
This figure visualizes the importance of directly targeting the intermediate layers to modify the representations within the model Jeon et al. (2024), where only targeting the output can remains the model unchanged internally.
B.5 Generation Results
In this section we show the output of different methods for harmful requests. Figure 8 displays cases where RepBend refuses to answer while other methods - NPO and CB - comply with a given request. RepBend refuses to answer by generating either a complete sentence of refusal or random tokens through the process visualized in Figure 3 and Figure 7.
B.6 Computing infrastructure
We utilized an on-demand computing environment of Yonsei University and Amazon AWS consisting of following Nvidia GPUs: A6000, A40, and L40S. Our initial exploratory research started with A6000 and A40 GPUs (around 500 hours considering single unit GPU usage). For hyper parameter search, we utilized sixteen L40S GPUs for 60 hours. For ablation and additional experiments during paper review period, we used 40 hours of sixteen L40S GPUs.